# Collaborative LLM Numerical Reasoning with Local Data Protection

### v1

@@ Abstract
Numerical reasoning over documents, which demands both contextual understanding and logical inference, is challenging for low-capacity local models deployed on computation-constrained devices. Although such complex reasoning queries could be routed to powerful remote models like GPT-4, exposing local data raises significant data leakage concerns. Existing mitigation methods generate problem descriptions or examples for remote assistance. However, the inherent complexity of numerical reasoning hinders the local model from generating logically equivalent queries and accurately inferring answers with remote guidance. In this paper, we present a model collaboration framework with two key innovations: (1) a context-aware synthesis strategy that shifts the query domains while preserving logical consistency; and (2) a tool-based answer reconstruction approach that reuses the remote-generated problem-solving pattern with code snippets. Experimental results demonstrate that our method achieves better reasoning accuracy than solely using local models while providing stronger data protection than fully relying on remote models. Furthermore, our method improves accuracy by 16.2% - 43.6% while reducing data leakage by 2.3% - 44.6% compared to existing data protection approaches.
1 Introduction
@@
Numerical reasoning over documents is a practical yet complex task that often requires powerful black-box models like GPT-4 for problem-solving (Akhtar et al., 2023). This task demands a deep understanding of documents, the ability to identify relationships from scattered evidence, and the capability to derive answers through quantitative calculations. In real-life scenarios, numerical reasoning is essential in tasks like analyzing financial reports (Chen et al., 2021; Zhao et al., 2022; Ma et al., 2025), research papers (Wu et al., 2025), medical documents (Mahendra et al., 2024), and contracts with numerical conditions (Huang et al., ). Due to the demanding requirements for contextual understanding and logical reasoning, on-device or in-house small models often struggle to solve such problems effectively. Consequently, remote black-box models with strong problem-solving capabilities are frequently accessed via API calls to address these challenges.
@@
However, the direct exposure of local data to remote models introduces significant risks of information leakage (Wang et al., 2024). Following prior studies (Zhou et al., 2023; Tong et al., 2023; Hartmann et al., 2024), we define local privacy information to pertain to every word, excluding non-sensitive stop words (Yue et al., 2021). Sensitive data can be presented explicitly or embedded implicitly within various contexts and formats, including company details, operational values, and strategic analyses, as illustrated in Fig. 1. While some works (Siyan et al., 2024; Chen et al., 2023c; Aahill, 2023) detect and remove explicit Personally Identifiable Information (PII), models like GPT-4 can infer personal attributes from residual context (Staab et al., ). Similarly, the sentence “our current policy is not to enter into transactions to hedge our fuel consumption…” in Fig. 1 reveals confidential policy decisions without containing any explicit PII. Therefore, in this work, we aim to thoroughly protect the local text before the black-box model inference.
Figure 1: (Left) An original example from the FinQA dataset. We highlight the sensitive information hidden in the context. (Right) The synthesis context and question with a different topic (from Aircraft to Healthcare) while preserving logical consistency, and corresponding problem-solving pattern from remote LLM for local recovery.
@@
The inherent complexity of numerical reasoning over documents exacerbates the trade-off between data leakage and model utility. On one hand, local data must be concealed from the remote model to minimize information leakage. On the other hand, the remote model often requires detailed contextual information to have a deep understanding of the problem and to provide accurate help for effective reasoning. We identify two primary challenges in protecting local data for numerical reasoning over documents:
@@
Difficulty in generating logically coherent synthesized queries. Existing methods often locally synthesize various queries for data protection, such as high-level descriptions (Zhang et al., 2024), analogous examples (Hartmann et al., 2024), dp-based permutation or paraphrased documents (Utpala et al., 2023). However, the complex logical reasoning within documents hampers the synthesis of accurate descriptions or logical-coherent contexts while keeping local data secret.
@@
Local answer reconstruction with limited reasoning ability. Existing methods integrate local data and the remote model’s response to generate or recover answers. Although they are effective for semantic-focused tasks such as classification (Utpala et al., 2023), translation, or summarization (Tong et al., 2023; Zhang et al., 2024; Hartmann et al., 2024), their performance drops dramatically in reasoning tasks which demands both deep contextual understanding and complex logic. Even with accurate hints or examples, small models often struggle to reconstruct correct answers due to their limited reasoning abilities.
To address these challenges, we propose a novel model collaboration framework based on a simply yet powerful intuition: we can transform sensitive reasoning requests into a different domain while maintaining their underlying logical structure. By translating reasoning requests from one domain (e.g., aircraft) to another (e.g., healthcare) (Fig. 1), our approach ensures that the remote model operates on semantically altered but logically equivalent data, thereby mitigating the risk of sensitive information exposure. Benefiting from the logic-preserving synthesis data, we design a tool-based strategy that elicits reusable problem-solving patterns from the remote model, facilitating accurate answer reconstruction. Our main contributions are summarized as follows:
  • We introduce a framework for collaborative numerical reasoning with black-box models, balancing local data protection and remote model utility without compromising local data. It seamlessly integrates with model cascade frameworks, leveraging remote models to compensate for local models’ limited reasoning ability.
  • We design a generalizable data protection approach using a distilled synthesis model to generate topic-shifted yet logic-consistent queries to seek help. Further, we propose a tool-based answer reconstruction method that leverages reusable problem-solving patterns from the remote model. Through these components, we achieve a superior accuracy-privacy trade-off than existing collaborative inference methods with data protection measures, improveing accuracy by 16.2% - 43.6% while reducing data leakage by 2.3% - 44.6%.
2 Background

### v2

@@ Abstract
Numerical reasoning over documents, which demands both contextual understanding and logical inference, is challenging for low-capacity local models deployed on computation-constrained devices. Although such complex reasoning queries could be routed to powerful remote models like GPT-4, exposing local data raises significant data leakage concerns. Existing mitigation methods generate problem descriptions or examples for remote assistance. However, the inherent complexity of numerical reasoning hinders the local model from generating logically equivalent queries and accurately inferring answers with remote guidance. In this paper, we present a model collaboration framework with two key innovations: (1) a context-aware synthesis strategy that shifts the query topics while preserving reasoning patterns; and (2) a tool-based answer reconstruction approach that reuses the remote-generated plug-and-play solution with code snippets. Experimental results demonstrate that our method achieves better reasoning accuracy than solely using local models while providing stronger data protection than fully relying on remote models. Furthermore, our method improves accuracy by 16.2% - 43.6% while reducing data leakage by 2.3% - 44.6% compared to existing data protection approaches.
Introduction
@@
Numerical reasoning over documents is a practical yet complex task that often requires powerful black-box models like GPT-4 for problem-solving (akhtar-etal-2023-exploring-nr). This task demands a deep understanding of documents, the ability to identify relationships from scattered evidence, and the capability to derive answers through quantitative calculations. In real-life scenarios, numerical reasoning is essential in tasks like analyzing financial reports (chen2021finqa; zhao2022multihiertt; ma2025llm_nr), research papers (wu2025agentic_nr), medical documents (mahendra-etal-2024-numbers_nr), and contracts with numerical conditions (huangnumerical_nr). Due to the demanding requirements for contextual understanding and logical reasoning, on-device or in-house small models often struggle to solve such problems effectively. Consequently, remote black-box models with strong problem-solving capabilities are frequently accessed via API calls to address these challenges.
@@
However, the direct exposure of local data to remote models introduces significant risks of information leakage (wang2024pandora). Following prior studies (zhou2023textobfuscator; tong2023inferdpt; hartmann2024can), we define local privacy information to pertain to every word, excluding non-sensitive stop words (yue2021differential). Sensitive data can be presented explicitly or embedded implicitly within various contexts and formats, including company details, operational values, and strategic analyses, as illustrated in Fig. 1 (a). While some works (siyan2024papillon; chen2023hide; Aahill2023AzureAI) detect and remove explicit Personally Identifiable Information (PII), models like GPT-4 can infer personal attributes from residual context (staabbeyond). Similarly, the sentence “our current policy is not to enter into transactions to hedge our fuel consumption…” in the example reveals confidential policy decisions without containing any explicit PII. Therefore, in this work, we aim to thoroughly protect the document-level local text before the black-box model inference.
Figure 1: (a) An example from FinQA with highlighted sensitive information. (b) The typical workflow and challenges for preserving privacy in interactions with remote API-based models.
@@
The inherent complexity of numerical reasoning over documents exacerbates the trade-off between data leakage and model utility. On one hand, local data must be concealed from the remote model to minimize information leakage. On the other hand, the remote model often requires detailed contextual information to have a deep understanding of the problem and to provide accurate help for effective reasoning. We identify two primary challenges for reasoning in the typical data protection workflow as shown in Fig. 1 (b):
@@
Difficulty in generating logically coherent synthesized queries. Existing methods often locally synthesize various queries for data protection, such as high-level descriptions (zhang2024cogenesis), analogous examples (hartmann2024can), dp-based permutation or paraphrased documents (utpala2023locally). However, the complex logical reasoning within documents hampers the synthesis of accurate descriptions or logical-coherent contexts while keeping local data secret.
@@
Local answer reconstruction with limited reasoning ability. Existing methods leverage a local model to integrate local data and the remote model’s response to generate final response. Although they are effective for semantic-focused tasks such as creative writing, translation and summarization (tong2023inferdpt; zhang2024cogenesis; hartmann2024can), their performance drops dramatically in reasoning tasks which demands both deep contextual understanding and complex logic. Even with accurate hints or examples, small models often struggle to reconstruct correct answers due to their limited reasoning abilities.
To address these challenges, we propose a novel method that transforms reasoning queries into a different domain while preserving the linguistic and logical patterns. As illustrated in (Fig. 2), by translating reasoning requests from one domain (e.g., aircraft fuel consumption) to another (e.g., advertising revenue) while preserving their inferential skeleton (e.g., deriving the total amount from the individual count and percentage), we enable secure delegation of reasoning to remote models without compromising sensitive information. This pattern-preserving transformation not only protects privacy but also retains the epistemic scaffolding necessary for solution generalization.
Building on this abstraction, we further introduce a tool-based answer reconstruction strategy where the remote model returns a reusable problem-solving tool with executable code snippets. This plug-and-play construct enables precise local answer recovery through direct numerical substitution, without local model dependence.
Our main contributions are summarized as follows:
  • We design a topic-shifted and pattern-preserving data synthesis approach that replaces the semantic surface while keeping the reasoning structure intact as a whole, enabling privacy-preserving delegation to remote models.
  • We propose a plug-and-play answer reconstruction method that leverages tool-centric reusable solution paradigm from the remote model, facilitating precise answer recovery via numerical substitution.
  • We achieve a superior accuracy-privacy trade-off than existing collaborative inference methods with data protection measures, improving accuracy by 16.2% - 43.6% while reducing data leakage by 2.3% - 44.6%.
Related Work
Prior works address data protection for training data (papernot2017semi; yue2021differential; tian2022seqpate; kurakin2023harnessing; xie2024differentially; yu2024privacy) or for demonstrations (hongdp; carey2024dp), but these approaches do not protect user queries during inference.
To protect privacy at inference time, some methods add differential privacy (DP) noise to text embeddings in white-box settings (du2023dp; zhou2023textobfuscator). However, these techniques are incompatible with black-box models, which typically require plain-text inputs rather than embeddings.
Some studies preserve privacy by replacing PII like names or locations using hider models (chen2023hide) or prompt optimization (siyan2024papillon), but residual context often still leaks information (staabbeyond).
For document-level privacy protection, some approaches apply DP-based perturbations to generate semantically similar but altered contexts (tong2023inferdpt; xie2024differentially), or prompt LLMs to paraphrase documents for downstream tasks like sentiment classification (utpala2023locally). Yet, these methods generally preserve semantics without guaranteeing logical consistency with the original content. Other works prompt local models to generate high-level descriptions (zhang2024cogenesis) or analogous examples (hartmann2024can). Still, generating queries that are both privacy-preserving and informative remains challenging, and local models often struggle to effectively utilize remote guidance due to limited capacity.
In contrast, our method enhances both privacy and logical fidelity by introducing a pattern-preserving topic shifter, and further employs a plug-and-play answer reconstruction strategy that avoiding local model dependence.
Method
Figure 2: The proposed method llustrated with examples for each step. The original query is transformed from one topic (e.g., aircraft fuel consumption) to another (e.g., advertising revenue) by a distilled topic rewriter, while preserving its reasoning pattern (e.g., deriving the total amount from the individual count and percentage). This enables secure delegation to a remote API-based model to elicit a tool-centric, plug-and-play solution for local answer reconstruction through direct numerical switch, without requiring local model re-inference.
To maximize the utility of remote models while minimizing local data leakage in numerical reasoning tasks, we propose an effective collaboration protocol based on the observation that sensitive queries can be transformed into a different domain without their underlying mathematical structure altered. Our approach involves translating queries from one domain to another, ensuring that the remote model processes semantically transformed but structurally equivalent data. This strategy mitigates the risk of exposing sensitive information while offering the local model a reusable solution. To achieve this, we introduce two key components of our protocol (as shown in Fig. 2), topic-shifted and pattern-preserving query synthesis and tool-based answer reconstruction, in the following subsections.
Topic-Shifted and Pattern-Preserving Query Synthesis
In this section, we introduce the module leveraging specialized local models to synthesize requests with shifted topics but equivalent mathematical abstractions. Since hiding the local information and maintaining the underlying logic remains challenging for small local models, we fine-tune a dedicated request synthesis model and subsequently apply a numerical replacement strategy. By decoupling semantic protection and numerical protection, maintaining numerical values intact in the topic shifter not only facilitates mathematical abstraction consistency verification but also paves the way for local answer reconstruction (as explained in the next section). Our data synthesis approach is detailed below.
Topic Shifter
To protect the overall local information instead of detecting specific sensitive words, we instruct the synthesis model to shift the topic while maintaining the original format, logic and numerical values. By preserving numeric values, we establish a clear mapping between the original and synthesized objects (e.g., fuel expense $\leftrightarrow$ advertising revenue in Fig. 2). We formally characterize the input and output of the proposed request synthesis model, denoted as $\mathcal{M_{S}}$, using the following equation:
$\quad\tilde{C},\tilde{q}=\mathcal{M_{S}}(C,q)$ (1)
where $\tilde{C},\tilde{q}$ represent the transformed context and query derived from the original inputs $C,q$.
To deploy an efficient local synthesizer, we distill the capabilities of a large remote model into a smaller local model. The remote model $\mathcal{M_{R}}$ generates synthetic data for training. Since the numerical values remain unchanged during the topic shift, a logically consistent rewritten question should yield the same answer as the original question. Through such post-hoc analysis, we confirm its proficiency in instruction-following and generating pattern-preserving requests. In contrast, small local models often struggle to meet these requirements, especially in keeping numerical consistency and logical coherence. Via instruction tuning, we enhance the local synthesizer’s rewriting and instruction-following capabilities.
Data Switch
Since the topic shifter focuses on semantic protection, leaving the numerical values unchanged, we further obfuscate numerical values in the synthesized request to ensure complete anonymization. Using regular expressions, we extract all numbers $\mathcal{N}=\{n_{1},n_{2},\dots,n_{k}\}$ from the request and apply a transformation
$h:n_{i}\mapsto\tilde{n_{i}}$ (2)
[…]
The data switch ensures numerical security, complementing the topic shifter’s for comprehensive data protection. Decoupling numerical protection during local data synthesis, we simplify subsequent local reconstruction. Serving as a bridge between the topic shifter and local reconstruction, the data switch ensures a seamless transition for accurate and secure data processing.
Tool-based Local Answer Reconstruction
Since the synthesized request maintains the same logic as the original one, they share the same problem-solving pattern. To best preserve and communicate this pattern, we instruct the remote model to generate Python code (chen2022program) as a tool for the local model. During answer reconstruction, the local model will simply perform data substitution: since intermediate steps are represented using Python variables, substituting the input data is sufficient. The final answer is obtained by executing the code in an interpreter. We elaborate each of these steps below.
Remote Assistance
After performing topic rewriting and numerical anonymization, we provide the synthesized context $\tilde{C_{h}}$ and synthesized question $\tilde{q_{h}}$ to the remote model $\mathcal{M_{R}}$:
[…]
where $m_{i}\in\{\tilde{n_{1}},\tilde{n_{2}},\dots,\tilde{n_{k}}\}$. The remote model returns the Python code snippets $f(\cdot)$ for problem solving with transformed numeric inputs from the synthesized request.
Local Reconstruction
Since the local model often fails to recognize the logical connection between the original and synthesized request and thus struggles to write Python codes for the local problem, we implement a plug-and-play approach to reuse the Python solution $f$ from the remote model. Specifically, we compute the answer to the local problem by directly executing the following with Python interpreter:
$f(h^{-1}(m_{1}),h^{-1}(m_{2}),\dots,h^{-1}(m_{t}))$ (4)
Recall that $h$ is the mapping between original and transformed numeric values. Simply by swapping the input values, we can obtain the answer for the local problem thanks to the logical consistency between original and synthesized requests. With the collaboration scheme above, we maximize the remote model utility while relieving the reasoning model from reasoning burdens. Before we move on to showcase the superior performance of our method, we would like to emphasize the unique contribution of our approach. While our method might seem to be an instantiation of Program-of-Thought (PoT)(chen2022program), our focus is mainly on how to reuse the Python solution with quite different contexts and accurately transfer to a logically equivalent problem instead of leveraging code to enhance reasoning. It’s important to note that PoT is merely a vehicle for our problem-solving pattern, and its success is deeply rooted in the logical consistency of queries as well as the decoupling of semantics and data protection within our approach.
Experiment Settings
Datasets MultiHiertt FinQA
[…]
Local Model: Phi-3-mini-128k-instruct (3.8B)
Local
Methods
Single-inference 54.4 0 71.1 0
Self-consistency 64.6 0 80.0 0
Collaboration
Methods
Hint 42.7 6.0 63.9 28.8
Example 57.0 16.1 71.4 38.9
Ours 80.1 3.7 87.6 6.4
Local Model: Llama-3.2-3B-Instruct
Local
Methods
Single-inference 33.8 0 54.0 0
Self-consistency 44.5 0 68.6 0
Collaboration
Methods
Hint 27.0 20.9 55.3 50.5
Example 31.1 18.8 55.5 20.0
Ours 70.6 5.5 90.1 5.9
Datasets
We conduct experiments using two question-answering datasets: FinQA (chen2021finqa) and MultiHiertt (zhao2022multihiertt). Both datasets involve questions that require numerical reasoning based on provided documents. They contain both explicit and implicit sensitive information, including financial data analysis, project details, and decision-making content, making them well-suited for evaluating our local data protection approach. Further details are provided in Appendix.
Data Processing
For privacy-preserving collaboration methods, we shorten the context by retrieving relevant parts of the full document. This retrieval process is a natural choice for long-document processing and reduces the data synthesis burden compared to using the entire document. To build it, we leverage the local self-consistency inference process by prompting the local model to surface supporting sentences during its reasoning. These are then combined with results from a BM25 retriever to form the final context. Details of the local retriever can be found in the Appendix, and it can also be replaced with other retrievers.
Models
We employed two lightweight local models designed for resource-constrained environments. Specifically, we used Phi-3-mini-128k-instruct (3.8B params), and Llama-3.2-3B-Instruct. For collaborative reasoning, we used GPT-4o as the remote model.
Model Distillation Settings
For the topic rewriter model, we adopted a distillation setup where Llama-3.2-3B-Instruct served as the student model and GPT-4o as the teacher model. After applying a data quality filtering process, via leakage evaluation, conflict evidence detection, and answer consistency verification (see Appendix), we retained 5762 training samples out of 6360 samples from MultiHiertt’s training set. We only trained a single synthesis model and used it for different datasets and local models.
Evaluation Metrics
Our evaluation comprises two key aspects: accuracy and local data leakage. Accuracy assesses whether the predicted answer matches the ground truth. Following (hartmann2024can), we report normalized accuracy scores (actual accuracy divided by remote-only inference accuracy) to show how much of the remote model’s performance different methods can achieve. For data leakage assessment, we follow prior studies (zhou2023textobfuscator; tong2023inferdpt), which define local privacy information as pertaining to every word, excluding non-sensitive stop words (yue2021differential). Similar to recent LLM-as-a-judge approaches (hartmann2024can; siyan2024papillon) for sensitive data leakage detection, we prompt the LLM to provide a binary judgment on whether the synthesized data contains original information. Specifically, we instruct a strong model to evaluate the presence of local information in remote model interactions. The evaluation prompt is as follows: Given context A and context B, determine whether context B uses information from context A. Ignore table formats and sentence structures; If they share some similar important nouns, it can be considered that context B uses information from context A. Respond directly with Yes or No. Here, context A denotes the original input, while context B represents the transmitted text during interactions with the remote model. We utilized GPT-4o-mini as the judge for data leakage evaluation. To validate the evaluation’s reliability, we conducted a manual review of 200 randomly selected leakage mapping results from the training set in both datasets. The agreement rate with human annotations was 96%, demonstrating a high alignment between the automated evaluation and human judgment.
Baselines
We compare our method with baselines involving local-only approaches, vanilla cascading without data protection, and model collaboration with different protection strategies. All methods utilize 3-shot prompting. We employ top-$p$ sampling (holtzmancurious) with $p=0.9$ for local models and greedy sampling for the remote model.
Single Inference: We use in-context learning with few-shot demonstrations (brown2020language) and Program-of-Thought (PoT) (chen2022program) for local solution generation. Results are averaged over seven runs.
Self-Consistency: To enhance local inference, we adopt self-consistency (wang2022self), selecting the answer with the highest voting consistency from seven executions.
Vanilla Model Cascading: Following (yuelarge), if local answer consistency falls below a threshold, the request is directly sent to a remote black-box model without data protection.
Hint: Following (zhang2024cogenesis; hartmann2024can), the local model generates a problem description, and the remote model provides a high-level hint. The hint is then integrated with local data for local inference.
Example: Following hartmann2024can; utpala2023locally, we use a local model to rephrase queries by concealing sensitive information (including both topics and numerical values) before sending to a remote model. The local model then combines local data with the remote solution for final inference.
Results
We first present the main results, comparing our method with local-only and existing privacy-preserving baselines in terms of accuracy and leakage relative to directly sending queries to the remote model. We then evaluate different privacy strategies under a model cascade framework, where a decision maker routes only harder queries to the remote model, to assess effectiveness on queries of different difficulty. Next, we conduct an ablation study to assess the contribution of each component. We further examine the impact of training data size on model performance. Finally, we perform an error analysis on incorrect cases.
Main Results
Table 1 shows the accuracy and leakage results for various methods with different local models on two datasets. Compared to local inference methods, our method provides an alternative solution that trades minimal data leakage for substantial accuracy improvements. Compared to collaborative inference methods with other data protection strategies, our method presents a superior improvement in both accuracy and data protection.
Our method achieves a new frontier in both the accuracy and the data protection over data protection baselines. (1) Accuracy improvement. Our method demonstrates a substantial performance improvement over existing approaches, with notable gaps observed across various datasets and models. Our method outperforms the Hint-based method by 23.7% - 43.6% in accuracy, and the Example-based method by 16.2% - 39.5%. Notably, the Hint and Example methods result in even lower accuracy than solely using the local model. This is because the unreliable, logic-incoherent query fails to trigger the remote large model for tailored assistance, and it does not effectively invoke the weaker local model’s reasoning capabilities. In contrast, our method generates logically consistent queries to elicit the remote model to produce problem-solving patterns, and it easily recovers answers through data replacement. (2) Leakage reduction. Our method achieves a significant reduction in data leakage compared to existing approaches. On the MultiHiertt dataset, leakage decreases by 2.3% - 15.4%. The reduction is even more pronounced on the FinQA dataset (14.1% - 44.6%). This is because the Hint method tends to leak contextual information when generating problem descriptions, and the Example method struggles to hide information effectively on its own. In contrast, our approach achieves effective information hiding and consistency maintenance by decoupling topic shifting and data replacement, along with distillation from the strong model. Examples of the synthesized query for different methods can be found in Appendix.
Our local synthesis model and overall framework are general, maintaining both model-agnostic and dataset-agnostic properties. Finally, our method demonstrates strong generalization across different datasets and local retrievers. Specifically, the local synthesis model was only trained on the MultiHiertt dataset using context shortened by the local retriever model Phi3. As shown in Table 1, our approach consistently achieves the highest accuracy and the lowest local data leakage rate, highlighting its effectiveness across unseen datasets (FinQA) and local retriever models (Llama3.2-3B-Instruct).
Results with a Referral Module in Model Cascade
In practice, it is unnecessary to forward all local instances to the remote model, as the local model can handle certain simple instances. Building on the vanilla model cascade (yuelarge), we use the local model’s answer consistency rate to determine whether the remote model should be involved. If the consistency rate is below the threshold, the remote model is activated to assist in the decision-making process. When the threshold is set to 0, all queries are handled locally; when set to 1, all queries involve the remote model.
Fig. 3 illustrates the accuracy and local data protection of various model collaboration methods at different thresholds. The solid lines represent accuracy, while the dashed lines indicate local data protection, with the local data protection rate calculated as 1 minus the leakage rate. The upper-right corner represents a combination of high accuracy and high local data protection. Our results show that while the vanilla model cascade method achieves higher accuracy by directly exposing the full original context, it leads to significant local data leakage. In contrast, baseline privacy-preserving methods improve local data protection but at the cost of dramatically reduced accuracy. Our approach strikes a balance, achieving comparable accuracy while maintaining high local data protection.
Ablation Study
Distil. Tool Acc Leak.
Ours 90.1 5.9
w/o Tool 65.0 5.9
w/o Distil. 42.2 15.1
w/o Distil. and Tool 65.7 15.1
Sensitivity to Training Data Size
Training Size Acc (%) Leakage (%)
800 74.3 9.2
2000 73.2 8.2
5762 70.6 5.5
Table 3: Sensitivity to training data size on MultiHiertt dataset with Llama3.2-3B-Instruct model.
We vary the training size for the topic rewriter and report the results in Table 3. As the training size decreases, we observe a slight increase in task accuracy accompanied by a consistent rise in leakage. This is because the topic rewriter sometimes makes only minor changes to the original query when trained on less data. Thus, the accuracy increases at the cost of reduced privacy protection. This aligns with the inherent accuracy–leakage trade-off. Remarkably, even with only 800 training examples, our method still significantly outperforms existing baselines in both accuracy and privacy, demonstrating the effectiveness of our approach.
Error Analysis
Error Type Percentage (%)
Retrieval Error 20
Rewrite Error 16.7
Numeric Switch Error 6.7
Remote LLM Error 33.3
Answer Reconstruction Error 0
Annotation Error/Ambiguous Question 23.3
Table 4: Error analysis.
We randomly sample 30 error cases from all incorrect predictions on MultiHiertt with the Phi3 model and manually analyze the error types. The distribution is shown in Table 4.
The most common source of error is remote LLM’s error (33.3%), including evidence grounding, logical reasoning, and Python coding errors. This is the imperfect model’s inherent error, note that the accuracy for remote-only method on MultiHiertt/FinQA dataset is 66.4% and 73.7%. A significant portion of the errors is attributed to flaws in the dataset itself (23.3%). Either the annotations were incorrect, or the questions were too ambiguous for a precise answer to be generated. The retrieval issues account for 20% of errors. We follow existing methods to shorten the context for long-text reasoning for input length-limited models as the first step in our method. The retrieval is not the main focus or contribution of our paper. It can be replaced with advanced retrieval methods to mitigate retrieval errors. The errors for two core designs in our method — topic rewriter (16.7%) and numeric switch (6.7%) for answer reconstruction (0%) are relatively small. Rewriter errors mainly stem from the rewriter model confusing some phrases that look similar or messing up the calculation units. Although imperfect, our design significantly reduces the errors compared to existing baselines by generating logically consistent synthesis data and reusing the Python tool with the data switch.
Conclusion
In this work, we propose a simple yet effective method for LLM numerical reasoning and protect the local data. We develop a context-aware synthesis strategy that ensures logical consistency while shifting query domains and a tool-based answer reconstruction approach that reuses remote-generated code snippets with local data. Our experimental results confirm the effectiveness of this approach, demonstrating substantial improvements in reasoning accuracy while mitigating data exposure. These findings pave the way for further advancements in privacy-preserving LLM collaboration systems, highlighting the potential for secure and efficient deployment in real-world applications.
Appendix A Appendix
Experiment Setting
This paper specifies the computing infrastructure used for running experiments (hardware and software), including GPU/CPU models; amount of memory; operating system; names and versions of relevant software libraries and frameworks
Why is the DP-based method unsuitable for our task?
The Differential Privacy (DP)-based method (tong2023inferdpt; xie2024differentially) for black-box inference adds noise to generate similar words as replacements for the original words. However, the resulting text is often unreadable and lacks logical coherence. While this approach may perform well on tasks like classification and sentiment analysis that primarily rely on semantic understanding, it is unsuitable for reasoning tasks that require logical consistency. An example generated using InferDPT (tong2023inferdpt) is provided below.
Original text:
He ’s been waiting 19 years for a visa still stuck in a backlog…
DP-generated text:
female declared billing 142 Pour must Fantasy Even dear poon am…
Model Cascade Framework